OpenStudy (anonymous):

how to integrate by substitution for (x^2+1)/(x+1) ?

OpenStudy (anonymous):

thank you for ur time, but im not taught using sin cos..

OpenStudy (anonymous):

\[ \frac{x^2+1}{x+1} = \frac{x^2-1+2}{x+1} = \frac{(x-1)(x+1)+2}{x+1}=x-1+\frac{2}{x+1} \]

OpenStudy (anonymous):

You can let \(u = x+1\)

OpenStudy (anonymous):

wio how do u get x2+1= x2-1+2?

OpenStudy (anonymous):

\(1 = 2-1 = -1 + 2\)

OpenStudy (anonymous):

hmm ok thx!

OpenStudy (anonymous):

why are both the answers so diff? im confused :(

OpenStudy (anonymous):

x²-x+2ln(x+1)

OpenStudy (anonymous):